OpenStudy (anonymous):

djano likes to roll the dice, but he hates to roll anything but 7 or 11. when other rolls come up he just ignores them and rolls again until he gets 7 or 11. if this is the case, what is the probability that he will roll a 7 before he rolls an 11?

OpenStudy (tkhunny):

What say you? First roll P(7) = 6/36 P(11) = 2/36 P(Anything Else) = 28/36 Did we get anywhere?

OpenStudy (anonymous):

where did the 6 and 2 come from